OpenStudy (anonymous):

why does ln9 x 0.5 = ln 3

OpenStudy (anonymous):

Because 0.5 * ln 9 = 0.5 * ln (3^2) By the rule \[\ln (b^c) = c*\ln (b)\] We get 0.5*2(ln3)=ln3

OpenStudy (anonymous):

@virtus , got it ?

OpenStudy (anonymous):

yes SIR!

OpenStudy (anonymous):

Thank you very much~

OpenStudy (anonymous):

LOL, don't call me sir :P and welcome :D

OpenStudy (anonymous):

i'll call you CAPTAIN then hahahah LOL jks